# HyperAdaLoRA: Accelerating LoRA Rank Allocation During Training via Hypernetworks without Sacrificing Performance

### v1

@@ 4.7 Efficiency Analysis
We analyze the computational load of our method compared to AdaLoRA. Table 7 presents the GPU memory usage and per step training latency for both methods under different batch sizes. As shown in Table 7, HyperAdaLoRA exhibits a slight reduction in memory usage compared to AdaLoRA. Additionally, HyperAdaLoRA consistently demonstrates lower training latency per step. Although the per step reduction may appear modest, the faster convergence rate demonstrated earlier leads to a significantly shorter total training time to reach the target performance level. Therefore, HyperAdaLoRA achieves a notable improvement in training efficiency.
This improvement in efficiency can be primarily attributed to two key factors. First, the hypernetwork we employ is lightweight and does not introduce significant additional computational overhead. Second, the latency reduction benefits from the convergence acceleration enabled by the hypernetwork. By generating task-specific parameters conditioned on contextual inputs, hypernetworks enable dynamic adaptation without requiring explicit iterative optimization, effectively capturing complex parameter adaptation patterns in real time. Hypernetworks facilitate adaptive parameter updates by modulating both the update direction and magnitude according to the current model state. This dynamic modulation promotes efficient exploration of high-dimensional parameter spaces, accelerating convergence while preserving model expressiveness.
@@ 5 Conclusion
In this paper, we tackle the issue of slow convergence in AdaLoRA, an effective dynamic rank allocation method for PEFT. We propose HyperAdaLoRA, a novel framework that leverages hypernetworks to dynamically generate the SVD-based parameters ($P,\Lambda,Q$) that are integral to AdaLoRA. By adopting an attention-based hypernetwork architecture, HyperAdaLoRA is capable of capturing intricate parameter dependencies and producing targeted updates. This enables it to navigate the optimization landscape more efficiently compared to traditional AdaLoRA training methods. Extensive experiments conducted on various datasets and models consistently show that HyperAdaLoRA achieves significantly faster convergence while maintaining performance. Additionally, further extension experiments on other LoRA-based approaches validate the broad applicability of our method.
@@ Appendix A Comprehensive Analysis of Hypernetworks for Accelerating Convergence
Appendix B Processing Details of Parameter Matrices (e.g., $P$)
@@
Appendix C Implementation Details of the Pruning Strategy
@@
which reflects both the magnitude of the singular value and its sensitivity to the gradient, indicating its contribution to parameter adaptation. After computing all scores, we identify the $k$ smallest $s_{ij}$ across all matrices and prune their associated singular values. This targeted pruning reduces the rank allocation for less significant directions and enables dynamic redistribution of the global rank budget, prioritizing more impactful components and improving overall parameter efficiency during training. Our parameter $k$ is fixed and predefined.
Appendix D Training Details
@@
During training, we employ a linear warm-up strategy to smoothly ramp up the learning rate, thereby enhancing training stability and convergence. Specifically, the learning rate is linearly increased from zero to the preset maximum value over the first 500 steps, after which the main learning rate scheduler (cosine annealing) takes over. Model parameters are initialized with a normal distribution. To ensure a fair comparison, all these training configurations are kept consistent.
Appendix E Case Study

### v2

@@ 4.7 Efficiency Analysis
We analyze the computational load of our method compared to AdaLoRA. Table 7 presents the GPU memory usage and per step training latency for both methods under different batch sizes. As shown in Table 7, HyperAdaLoRA exhibits a slight reduction in memory usage compared to AdaLoRA. Additionally, HyperAdaLoRA consistently demonstrates lower training latency per step. Although the per step reduction may appear modest, the faster convergence rate demonstrated earlier leads to a significantly shorter total training time to reach the target performance level. Therefore, HyperAdaLoRA achieves a notable improvement in training efficiency. This efficiency gain mainly stems from two factors: first, our hypernetwork is lightweight and adds minimal computational overhead; second, by generating task-specific parameters conditioned on context, it enables dynamic adaptation without iterative optimization, accelerating convergence. By modulating update direction and magnitude, it efficiently explores high-dimensional parameter spaces while preserving model expressiveness. We provide an analysis of memory usage in Appendix I.
@@ 5 Conclusion
In this paper, we address the issue of slow convergence in AdaLoRA, an effective dynamic rank allocation method within the PEFT framework. To overcome this limitation, we propose HyperAdaLoRA, a novel and flexible framework that employs hypernetworks to dynamically generate the SVD-based low-rank parameters $(P,\Lambda,Q)$ central to AdaLoRA. By adopting an attention-enhanced hypernetwork architecture, HyperAdaLoRA captures fine-grained parameter dependencies and produces more targeted updates during training. This design enables the model to explore the optimization space more efficiently, leading to notably faster convergence compared to standard AdaLoRA training methods. Extensive experiments across multiple benchmarks and model scales show that HyperAdaLoRA achieves faster convergence while maintaining comparable or even better final performance. Furthermore, extension experiments on other LoRA-based methods clearly demonstrate the versatility and broad applicability of our approach.
@@ Appendix A Comprehensive Analysis of Hypernetworks for Accelerating Convergence
Appendix B Details of Hypernetworks for LoRA, DoRA, and DyLoRA
In line with our formulation that extends AdaLoRA to HyperAdaLoRA in the paper, we now provide a parallel description of how hypernetworks are applied to LoRA, DoRA, and DyLoRA. For LoRA and DyLoRA, their weight update formula is the same as in standard LoRA: $\Delta W=BA$, where $B$ and $A$ are updated via backpropagation. We introduce two hypernetworks, $H_{B}$ and $H_{A}$, to generate the updated parameters: $B_{i+1}=H_{B}(B_{i};\Phi_{B})$, $A_{i+1}=H_{A}(A_{i};\Phi_{A})$, where $\Phi_{B}$ and $\Phi_{A}$ are the trainable parameters of the hypernetworks. $B_{i}$ and $A_{i}$ denote the parameters before the $i$-th iteration, and $B_{i+1}$ and $A_{i+1}$ denote the parameters after the update. For DoRA, the original method factorizes the LoRA low-rank update into a direction component and a scalar component. Let the directional part be $\Delta W_{\mathrm{dir}}=B_{\mathrm{dir}}A_{\mathrm{dir}}$, and let the scalar be $s$. The weight update can then be written as $\Delta W=s\cdot\Delta W_{\mathrm{dir}}=s\cdot B_{\mathrm{dir}}A_{\mathrm{dir}}$. In standard DoRA, $B_{\mathrm{dir}}$ and $A_{\mathrm{dir}}$ are directly updated via backpropagation. Analogously, we introduce hypernetworks for the directional component in DoRA. In particular, we use two hypernetworks, $H_{B_{\mathrm{dir}}}$ and $H_{A_{\mathrm{dir}}}$, to generate the updated parameters: $B_{\mathrm{dir},i+1}=H_{B_{\mathrm{dir}}}(B_{\mathrm{dir},i};\Phi_{B_{\mathrm{dir}}})$, $A_{\mathrm{dir},i+1}=H_{A_{\mathrm{dir}}}(A_{\mathrm{dir},i};\Phi_{A_{\mathrm{dir}}})$, where $\Phi_{B_{\mathrm{dir}}}$ and $\Phi_{A_{\mathrm{dir}}}$ are the trainable parameters of the hypernetworks.
Appendix C Processing Details of Parameter Matrices (e.g., $P$)
@@ Appendix B Details of Hypernetworks for LoRA, DoRA, and DyLoRA
Appendix D Implementation Details of the Pruning Strategy
@@
which reflects both the magnitude of the singular value and its sensitivity to the gradient, indicating its contribution to parameter adaptation. After computing all scores, we identify the $k$ smallest $s_{ij}$ across all matrices and prune their associated singular values. This targeted pruning reduces the rank allocation for less significant directions and enables dynamic redistribution of the global rank budget, prioritizing more impactful components and improving overall parameter efficiency during training. Our parameter $k$ is fixed and predefined. It is worth noting that the pruning strategy is not the primary focus or core contribution of our work. It is directly inherited from the original design of AdaLoRA and is used solely to maintain consistency and fairness with prior work, rather than to introduce a new pruning technique.
Appendix E Other Training Details
@@
During training, we employ a linear warm-up strategy to smoothly ramp up the learning rate, thereby enhancing training stability and convergence. Specifically, the learning rate is linearly increased from zero to the preset maximum value over the first 500 steps, after which the main learning rate scheduler (cosine annealing) takes over. Model parameters are initialized with a normal distribution. We set the average budget per weight to three, giving a total budget equal to three times the number of weights, and use a pruning interval of 100 steps. To ensure a fair comparison, all these training configurations are kept consistent.
Appendix F $P\Lambda Q$ Parameter Similarity Analysis
We conduct an average similarity analysis between the $(P\Lambda Q)$ parameters generated by our method and those produced by AdaLoRA. As shown in Table 8, the results indicate that the similarity between corresponding parameters consistently exceeds 0.98, demonstrating that the two methods are highly consistent at the parameter level and essentially equivalent in terms of performance. This finding further supports that our method can achieve comparable parameter representations while maintaining strong performance.
Parameter $P$ $\Lambda$ $Q$
Cosine Similarity 0.9861 0.9878 0.9811
Table 8: Cosine similarity between the $(P\Lambda Q)$ parameters of our method and AdaLoRA.
Appendix G Additional Results on Training Time and Performance
We further present a comparative performance evaluation on the OpenPlatypus dataset, as well as a training time comparison of LLaMA3.1-8B on GSM8K and HumanEval. The corresponding results are summarized in Table 9 and 10. These results demonstrate that our method improves training efficiency without compromising performance.
Model Method BLEU-4 ROUGE-1
Qwen2.5-7B AdaLoRA 19.87 44.24
Ours 19.92 44.65
LLaMA3.1-8B AdaLoRA 34.86 52.90
Ours 35.95 53.00
Table 9: Performance comparison of HyperAdaLoRA and AdaLoRA on OpenPlatypus.
Dataset LoRA w/o LoRA w/ AdaLoRA w/o AdaLoRA w/ DoRA w/o DoRA w/ DyLoRA w/o DyLoRA w/
GSM8K 9863 8110 10542 9011 10892 9054 12305 10527
HumanEval 8625 7210 10281 8620 9532 8029 10085 8547
Table 10: Comparison of training time (s) across methods with and without hypernetworks (w/o = without hypernetworks, w/ = with hypernetworks).
Appendix H Comparison with the $\gamma$ = 0 Setting (without regularization)
We evaluate our method under the $\gamma=0$ setting on the LLaMA3.1-8B model using the Stanford Alpaca and Magpie datasets. As shown in Table 11, setting $\gamma=0$ leads to a slight, yet acceptable, drop in accuracy, while the training time remains largely unchanged. This indicates that our method can still achieve strong performance even without applying regularization.
Method Stanford Alpaca Magpie Stanford Alpaca (Time) Magpie (Time)
Ours ($\gamma=0$) 54.88 70.36 6970 14920
Ours 55.10 70.73 6650 15720
Table 11: Evaluation of our method under $\gamma=0$ on the Stanford Alpaca and Magpie datasets.
Appendix I Memory Analysis
We provide here an analysis of how our method reduces memory consumption. The number of parameters introduced by our hypernetwork is much smaller than the total number of fine-tuning parameters introduced in the baselines (e.g., $(P,\Lambda,Q)$ in AdaLoRA, or $(A,B)$ in LoRA). In these baseline methods, the optimizer needs to maintain separate first- and second-moment states for all these fine-tuning parameters, which consumes a considerable amount of memory. In our method, $(P,\Lambda,Q)$ / $(A,B)$ are generated on demand by a shared lightweight hypernetwork and are no longer treated as independent trainable parameters, so the optimizer only needs to maintain states for the hypernetwork parameters. This substantially reduces the amount of optimizer states, leading to a slightly lower peak memory footprint.
Appendix J Training Stability of Our Method
Our proposed architecture is inherently stable to train for several reasons: (1) Shared optimization objective: the hypernetwork does not alter the main model’s optimization objective but merely imposes a structured parametrization on the LoRA parameters, so the system still minimizes the same task loss instead of engaging in an adversarial game; (2) Low-dimensional update space: the hypernetwork only operates on low-rank LoRA factors, which imposes strong constraints on the parameter space and restricts updates to a structured low-dimensional subspace, leading to a smoother and more tractable optimization landscape; (3) Aligned gradient signals: the main model and the hypernetwork share the same task loss and gradient signal, so there are no competing objectives and their optimization directions are statistically aligned; (4) Lightweight hypernetwork design: the hypernetwork is relatively small and trained with conservative optimization hyperparameters, behaving more like a lightweight adaptation module rather than a second large backbone; (5) LoRA bounded update magnitude: the low-rank structure of LoRA naturally bounds the magnitude of updates and gradient norms, so even when local gradients are large, the effective update space is constrained by the rank and bottleneck dimensions, which in practice helps prevent gradient explosion and numerical divergence.
Appendix K Case Study